# A Survey of Reinforcement Learning For Economics

### 23 Mar 2026 v4

@@ 5.5.2 Entropy Regularization and Soft Actor-Critic
where $\mathcal{H}(\pi)=-\sum_{a}\pi(a)\log\pi(a)$ is the entropy and $\tau>0$ is the temperature parameter. Geist et al. (2019) later provided the unifying theoretical framework, showing that entropy regularization converts the Bellman optimality operator’s non-smooth hard max into a smooth log-sum-exp, and that the resulting soft Bellman operator remains a $\gamma$-contraction, preserving the convergence guarantees of standard dynamic programming.102102102The optimal policy under entropy regularization is $\pi^{*}(a|s)\propto\exp(Q^{*}(s,a)/\tau)$, which is precisely the McFadden (1974) logit choice probability with systematic utility $Q^{*}(s,a)$ and scale parameter $\tau$. The entropy-regularized value function is the log-sum-exp of Q-values, corresponding to the inclusive value (log-sum) operator in nested logit models. This equivalence between the soft-control framework and dynamic discrete choice models with EV1 taste shocks is developed formally in Rust and Rawat (2026), Appendix A.
@@ 5.5.3 Error Amplification Under Approximate Value Functions
At $\gamma=0.99$, the amplification factor is $2\cdot 0.99/0.01=198$. A 1% error in value function approximation yields at most 198% error in policy value.103103103The Singh-Yee bound is worst-case and not tight in general; massoud Farahmand et al. (2010) derive tighter bounds under smoothness assumptions on the MDP. The amplification factor $2\gamma/(1-\gamma)$ is a sensitivity analysis: it quantifies how errors in the “inputs” (value estimates) propagate to “outputs” (policy quality), analogous to errors-in-variables bias in regression. The discount factor $\gamma$ controls sensitivity; more patient agents face larger amplification. This bound is pessimistic but finite: approximate value functions do not cause unbounded policy degradation.
@@ 5.5.4 Sample Complexity of Planning
Classical dynamic programming complexity scales with the state space size $|\mathcal{S}|$. For problems like Go, where $|\mathcal{S}|\approx 10^{170}$, exact computation is impossible. Kearns et al. (2002) prove that with access to a generative model104104104A “generative model” in RL is a simulator that, given any state-action pair $(s,a)$, returns a sampled next state $s^{\prime}\sim P(\cdot|s,a)$ and reward $r$. This is unrelated to “generative models” in machine learning (GANs, diffusion models) or “generative processes” in Bayesian econometrics. The distinction matters: planning with a generative model is strictly easier than learning from a single trajectory, because the agent can query arbitrary states rather than following a sequential path. (a simulator that samples transitions from any state-action pair), near-optimal planning is possible with no dependence on $|\mathcal{S}|$. The cost is exponential dependence on the effective horizon $H=\log(R_{\max}/(\epsilon(1-\gamma)))/\log(1/\gamma)$: the sparse sampling algorithm requires $O((|\mathcal{A}|/\epsilon)^{H})$ simulator calls.105105105The sparse sampling algorithm builds a random tree of depth $H$ from the current state, sampling $C$ successor states per action at each node, then estimates values by averaging leaf rewards back up the tree. Its running time is $O((C|\mathcal{A}|)^{H})$, which is exponential in $H$ but entirely independent of $|\mathcal{S}|$. Kearns et al. (2002) also prove a lower bound of $\Omega(2^{H})$ generative model calls for any planning algorithm, so exponential horizon dependence is unavoidable in the worst case. For $\gamma$ near 1, $H\approx(1-\gamma)^{-1}\log(R_{\max}/\epsilon)$, so the method is practical only for short effective horizons or moderate discount factors. The key insight is the tradeoff: classical DP scales linearly in $|\mathcal{S}|$ but polynomially in $H$; sparse sampling eliminates state-space dependence at the cost of exponential horizon dependence. This explains why MCTS succeeds in large state spaces with bounded lookahead.
@@
The minimax-optimal sample complexity for planning with a generative model, when queries to arbitrary state-action pairs are permitted, is $\Theta(|\mathcal{S}||\mathcal{A}|/((1-\gamma)^{3}\epsilon^{2}))$ (Azar et al., 2013). This bound scales linearly in $|\mathcal{S}|$ but polynomially in $1/(1-\gamma)$, the opposite regime from sparse sampling. Agarwal et al. (2020a) show that the plug-in model-based approach (learn $\hat{P}$ from samples, then plan with $\hat{P}$) achieves this minimax rate, establishing that model-based RL is statistically optimal. Li et al. (2024b) further tighten this result by breaking the $|\mathcal{S}||\mathcal{A}|/(1-\gamma)^{2}$ sample-size barrier, showing that the minimax rate is achievable with total sample size as low as $|\mathcal{S}||\mathcal{A}|/(1-\gamma)$.106106106Recent extensions push these results beyond standard MDPs. Clavier et al. (2024) study the robust MDP setting where the agent must plan under model uncertainty with sa-rectangular or s-rectangular uncertainty sets, establishing minimax rates for robust policy optimization. Wang et al. (2025) extend the analysis to risk-sensitive objectives under the iterated CVaR criterion.
@@ 5.6 Fundamental Tradeoffs
Four additional trade-offs pervade reinforcement learning. First, exploration versus exploitation: should the agent act on its current best estimate or gather information to improve future decisions? Lai and Robbins (1985) establish the fundamental lower bound: any consistent policy must incur regret at least logarithmic in the number of periods. Naive exploration ($\varepsilon$-greedy) requires samples exponential in the horizon; strategic exploration (UCB, optimism in the face of uncertainty) reduces this to polynomial (Auer et al., 2002a), formalizing the value of targeted experimentation.107107107The exploration-exploitation tradeoff is the subject of the bandits chapter, where the multi-armed bandit framework provides the sharpest analysis. In full MDPs, exploration is harder because the agent must learn not just reward distributions but also transition dynamics, compounding the information requirement.
Second, model-based versus model-free: model-based methods learn a transition model $\hat{P}(s^{\prime}|s,a)$ and plan with it (Sutton, 1990); model-free methods learn value functions or policies directly from transitions. The Dyna architecture (Sutton, 1990) bridges these by generating simulated experience from the learned model to supplement real transitions. Model-based methods are sample-efficient (each transition updates the entire model, which improves value estimates for all states) but suffer asymptotic bias if the model class is misspecified; model-free methods are asymptotically unbiased but sample-inefficient, using each transition for a single gradient step.108108108Model misspecification in RL is the analog of omitted variable bias in econometrics: if the learned model omits relevant state variables or misspecifies the functional form of transitions, the resulting policy is biased regardless of sample size. Moerland et al. (2023) provide a comprehensive survey of model-based RL, analyzing the model-bias versus sample-efficiency tradeoff across method families.
@@ 6.1 The Moving Target Problem
Fujimoto et al. (2022) formalize this observation. Let $Q^{\pi}$ denote the true value function for policy $\pi$, let $\Delta(s,a)=Q(s,a)-Q^{\pi}(s,a)$ denote the value error, and let $\varepsilon(s,a)=Q(s,a)-(r+\gamma\mathbb{E}_{s^{\prime},a^{\prime}}[Q(s^{\prime},a^{\prime})])$ denote the Bellman error. Substituting the definition of $Q^{\pi}$ yields the key identity: $\varepsilon(s,a)=\Delta(s,a)-\gamma\mathbb{E}_{s^{\prime},a^{\prime}}[\Delta(s^{\prime},a^{\prime})]$. The Bellman error is a difference of value errors at consecutive states, not the value error itself. If the errors $\Delta(s,a)$ and $\Delta(s^{\prime},a^{\prime})$ are correlated across time—the network is wrong in the same direction at successive state-action pairs—they cancel in the difference and the Bellman error is small regardless of how large the individual errors are.109109109In the extreme case, shifting all Q-values by the constant $c/(1-\gamma)$ leaves the Bellman error unchanged at zero while increasing value error by $c/(1-\gamma)$.
The second failure mode is specific to finite datasets: Fujimoto et al. (2022, Corollary 1) show that over an incomplete dataset, zero Bellman error is consistent with arbitrarily large value error, because the network can fit unobserved successor pairs to whatever values make the observed residuals vanish.110110110The Bellman equation uniquely identifies $Q^{\pi}$ when enforced over the entire MDP, but over an incomplete dataset it admits infinitely many solutions. Whenever a transition $(s^{\prime},a^{\prime})$ that is reachable from the dataset is not itself in the dataset, the network is free to set $Q(s^{\prime},a^{\prime})$ at unobserved pairs to whatever value makes the residual vanish on the observed ones, unconstrained by any loss. A network can thus reach near-zero training loss while the value function remains arbitrarily inaccurate over the full state space.
The dual failure mode appears in policy gradient methods: Ilyas et al. (2020) find that even when PPO’s surrogate objective improves monotonically, episode return can plateau or decline, because the surrogate gradient is poorly aligned with the gradient of the true return.111111111Ilyas et al. (2020) examine the surrogate objective used in Proximal Policy Optimization (Schulman et al., 2017). The PPO clipping mechanism is designed to keep policy updates within a trust region by bounding the probability ratio $\pi_{\theta}(a|s)/\pi_{\theta_{\text{old}}}(a|s)$. The gradient of the surrogate is poorly aligned with the gradient of the true return, particularly in later training phases where the policy has diverged from the behavior policy used to collect the replay data. The loss metric that practitioners monitor throughout training is measuring something other than what they care about.
@@ 6.2 The Reproducibility Crisis and Sensitivity to Random Seeds
Henderson et al. (2018) trained five leading policy gradient algorithms (PPO, TRPO, DDPG, TD3, SAC) on six MuJoCo benchmark environments, holding all hyperparameters fixed and varying only the random seed. The resulting learning curves from different seeds were non-overlapping: a seed that performed well under one algorithm performed comparably to a different algorithm’s best seeds, making cross-algorithm comparison unreliable.112112112Differences as large as 2,000 points in final episode return arose from seed variation alone. Agarwal et al. (2021b) quantify the damage: comparing point estimates from 5 runs per task on Atari 100k yields Type I error exceeding 50%, meaning a random noise injection appears beneficial in half of all comparisons.
Agarwal et al. (2021b) propose the interquartile mean (IQM) as a replacement for mean and median when comparing algorithms. The IQM discards the top and bottom 25% of runs before averaging, reducing sensitivity to outlier seeds. Using these tools and stratified bootstrap confidence intervals, Agarwal et al. (2021b) find that several widely-cited algorithmic improvements on Atari 100k vanish or reverse when statistical uncertainty is accounted for.113113113They also introduce performance profiles, which plot the fraction of tasks and seeds where an algorithm achieves performance above a threshold $\tau$, as $\tau$ varies from 0 to the maximum. Performance profiles reveal the full shape of the score distribution rather than collapsing it to a single statistic.114114114The fragility extends to hyperparameters. Eimer et al. (2023) conduct a systematic study of hyperparameter sensitivity across 6 algorithms and 17 environments, finding that default hyperparameters from published papers perform competitively in the specific environments used in those papers but generalize poorly across environments. Patterson et al. (2024) synthesize these findings into an empirical design handbook, recommending at minimum 10 seeds per configuration, IQM-based comparisons, and preregistration of hyperparameter search protocols.
@@ 6.3 Value Overestimation and Spikes
Fujimoto et al. (2018) observe that Double DQN’s correction is incomplete in continuous-action settings, where the target network and online network remain correlated through shared updates. They propose Clipped Double Q-learning: compute two Q-value estimates $Q_{1},Q_{2}$ with separate networks trained on the same data, and use $y=r+\gamma\min(Q_{1}(s^{\prime},a^{\prime}),Q_{2}(s^{\prime},a^{\prime}))$ as the Bellman target. The minimum operator introduces pessimistic underestimation, which is conservative but avoids the explosive positive bias.115115115Ciosek et al. (2019) note that clipped double Q can cause excessive pessimism under high uncertainty, proposing optimistic actor-critic as a counterweight.
@@
DQN prevents outright divergence, but van Hasselt et al. (2018) find that soft divergence, defined as a temporary spike in value estimates by more than 10% followed by recovery, occurs in the majority of DQN training runs across 57 Atari games.116116116Larger networks diverge more frequently than smaller ones, counter to the usual intuition that more expressive models should generalize better. The deadly triad does not announce itself as a training failure: the value estimates may spike and recover, leaving no visible trace in the loss curve while corrupting the policy.
Kumar et al. (2021) describe a continuous manifestation of the deadly triad: implicit under-parameterization.117117117Neural networks trained with bootstrapped TD objectives progressively lose their effective rank, with fewer and fewer neurons contributing distinct directions in the representation. This rank collapse is silent (training loss continues to decrease) but the network’s ability to represent new information degrades over time, a form of capacity loss distinct from but related to plasticity loss.
@@ 6.4 Plasticity Loss and Primacy Bias
Nikishin et al. (2022) identify primacy bias as a specific cause. Because the replay buffer is filled incrementally, early transitions are oversampled relative to later ones, and the network over-fits to early environment transitions, corrupting representations throughout the remainder of training.118118118Nikishin et al. (2022) show that 100 initial “priming” steps of excessive gradient updates degrade a SAC agent’s performance for hundreds of thousands of subsequent steps. Sokar et al. (2023) measure the fraction of dormant neurons (units with near-zero activation across the replay buffer) accumulating monotonically during training. Dohare et al. (2024) find that standard deep networks lose all plasticity within a few million gradient steps in continual learning tasks.
The proposed remedies fall into three categories: periodic resets, continual backpropagation, and architectural interventions.119119119Periodic resets reinitialize the last layers of the network while retaining the replay buffer, allowing the agent to forget overfit representations without discarding experience (Nikishin et al., 2022; D’Oro et al., 2023). Continual backpropagation replaces neurons with near-zero utility at each gradient step rather than waiting for a global reset (Dohare et al., 2024). Architectural interventions—layer normalization (Lyle et al., 2025), orthogonal initialization, spectral normalization—reduce the rate at which plasticity is lost by stabilizing gradient magnitudes and preventing weight norm growth.

### 24 Mar 2026 v5

@@ 5.5.2 Entropy Regularization and Soft Actor-Critic
where $\mathcal{H}(\pi)=-\sum_{a}\pi(a)\log\pi(a)$ is the entropy and $\tau>0$ is the temperature parameter. Geist et al. (2019) later provided the unifying theoretical framework, showing that entropy regularization converts the Bellman optimality operator’s non-smooth hard max into a smooth log-sum-exp, and that the resulting soft Bellman operator remains a $\gamma$-contraction, preserving the convergence guarantees of standard dynamic programming.104104104The optimal policy under entropy regularization is $\pi^{*}(a|s)\propto\exp(Q^{*}(s,a)/\tau)$, which is precisely the McFadden (1974) logit choice probability with systematic utility $Q^{*}(s,a)$ and scale parameter $\tau$. The entropy-regularized value function is the log-sum-exp of Q-values, corresponding to the inclusive value (log-sum) operator in nested logit models. This equivalence between the soft-control framework and dynamic discrete choice models with EV1 taste shocks is developed formally in Rust and Rawat (2026), Appendix A.
@@ 5.5.3 Error Amplification Under Approximate Value Functions
At $\gamma=0.99$, the amplification factor is $2\cdot 0.99/0.01=198$. A 1% error in value function approximation yields at most 198% error in policy value.105105105The Singh-Yee bound is worst-case and not tight in general; massoud Farahmand et al. (2010) derive tighter bounds under smoothness assumptions on the MDP. The amplification factor $2\gamma/(1-\gamma)$ is a sensitivity analysis: it quantifies how errors in the “inputs” (value estimates) propagate to “outputs” (policy quality), analogous to errors-in-variables bias in regression. The discount factor $\gamma$ controls sensitivity; more patient agents face larger amplification. This bound is pessimistic but finite: approximate value functions do not cause unbounded policy degradation.
@@ 5.5.4 Sample Complexity of Planning
Classical dynamic programming complexity scales with the state space size $|\mathcal{S}|$. For problems like Go, where $|\mathcal{S}|\approx 10^{170}$, exact computation is impossible. Kearns et al. (2002) prove that with access to a generative model106106106A “generative model” in RL is a simulator that, given any state-action pair $(s,a)$, returns a sampled next state $s^{\prime}\sim P(\cdot|s,a)$ and reward $r$. This is unrelated to “generative models” in machine learning (GANs, diffusion models) or “generative processes” in Bayesian econometrics. The distinction matters: planning with a generative model is strictly easier than learning from a single trajectory, because the agent can query arbitrary states rather than following a sequential path. (a simulator that samples transitions from any state-action pair), near-optimal planning is possible with no dependence on $|\mathcal{S}|$. The cost is exponential dependence on the effective horizon $H=\log(R_{\max}/(\epsilon(1-\gamma)))/\log(1/\gamma)$: the sparse sampling algorithm requires $O((|\mathcal{A}|/\epsilon)^{H})$ simulator calls.107107107The sparse sampling algorithm builds a random tree of depth $H$ from the current state, sampling $C$ successor states per action at each node, then estimates values by averaging leaf rewards back up the tree. Its running time is $O((C|\mathcal{A}|)^{H})$, which is exponential in $H$ but entirely independent of $|\mathcal{S}|$. Kearns et al. (2002) also prove a lower bound of $\Omega(2^{H})$ generative model calls for any planning algorithm, so exponential horizon dependence is unavoidable in the worst case. For $\gamma$ near 1, $H\approx(1-\gamma)^{-1}\log(R_{\max}/\epsilon)$, so the method is practical only for short effective horizons or moderate discount factors. The key insight is the tradeoff: classical DP scales linearly in $|\mathcal{S}|$ but polynomially in $H$; sparse sampling eliminates state-space dependence at the cost of exponential horizon dependence. This explains why MCTS succeeds in large state spaces with bounded lookahead.
@@
The minimax-optimal sample complexity for planning with a generative model, when queries to arbitrary state-action pairs are permitted, is $\Theta(|\mathcal{S}||\mathcal{A}|/((1-\gamma)^{3}\epsilon^{2}))$ (Azar et al., 2013). This bound scales linearly in $|\mathcal{S}|$ but polynomially in $1/(1-\gamma)$, the opposite regime from sparse sampling. Agarwal et al. (2020a) show that the plug-in model-based approach (learn $\hat{P}$ from samples, then plan with $\hat{P}$) achieves this minimax rate, establishing that model-based RL is statistically optimal. Li et al. (2024b) further tighten this result by breaking the $|\mathcal{S}||\mathcal{A}|/(1-\gamma)^{2}$ sample-size barrier, showing that the minimax rate is achievable with total sample size as low as $|\mathcal{S}||\mathcal{A}|/(1-\gamma)$.108108108Recent extensions push these results beyond standard MDPs. Clavier et al. (2024) study the robust MDP setting where the agent must plan under model uncertainty with sa-rectangular or s-rectangular uncertainty sets, establishing minimax rates for robust policy optimization. Wang et al. (2025) extend the analysis to risk-sensitive objectives under the iterated CVaR criterion.
@@ 5.6 Fundamental Tradeoffs
Four additional trade-offs pervade reinforcement learning. First, exploration versus exploitation: should the agent act on its current best estimate or gather information to improve future decisions? Lai and Robbins (1985) establish the fundamental lower bound: any consistent policy must incur regret at least logarithmic in the number of periods. Naive exploration ($\varepsilon$-greedy) requires samples exponential in the horizon; strategic exploration (UCB, optimism in the face of uncertainty) reduces this to polynomial (Auer et al., 2002a), formalizing the value of targeted experimentation.109109109The exploration-exploitation tradeoff is the subject of the bandits chapter, where the multi-armed bandit framework provides the sharpest analysis. In full MDPs, exploration is harder because the agent must learn not just reward distributions but also transition dynamics, compounding the information requirement.
Second, model-based versus model-free: model-based methods learn a transition model $\hat{P}(s^{\prime}|s,a)$ and plan with it (Sutton, 1990); model-free methods learn value functions or policies directly from transitions. The Dyna architecture (Sutton, 1990) bridges these by generating simulated experience from the learned model to supplement real transitions. Model-based methods are sample-efficient (each transition updates the entire model, which improves value estimates for all states) but suffer asymptotic bias if the model class is misspecified; model-free methods are asymptotically unbiased but sample-inefficient, using each transition for a single gradient step.110110110Model misspecification in RL is the analog of omitted variable bias in econometrics: if the learned model omits relevant state variables or misspecifies the functional form of transitions, the resulting policy is biased regardless of sample size. Moerland et al. (2023) provide a comprehensive survey of model-based RL, analyzing the model-bias versus sample-efficiency tradeoff across method families.
@@ 6.1 The Moving Target Problem
Fujimoto et al. (2022) formalize this observation. Let $Q^{\pi}$ denote the true value function for policy $\pi$, let $\Delta(s,a)=Q(s,a)-Q^{\pi}(s,a)$ denote the value error, and let $\varepsilon(s,a)=Q(s,a)-(r+\gamma\mathbb{E}_{s^{\prime},a^{\prime}}[Q(s^{\prime},a^{\prime})])$ denote the Bellman error. Substituting the definition of $Q^{\pi}$ yields the key identity: $\varepsilon(s,a)=\Delta(s,a)-\gamma\mathbb{E}_{s^{\prime},a^{\prime}}[\Delta(s^{\prime},a^{\prime})]$. The Bellman error is a difference of value errors at consecutive states, not the value error itself. If the errors $\Delta(s,a)$ and $\Delta(s^{\prime},a^{\prime})$ are correlated across time—the network is wrong in the same direction at successive state-action pairs—they cancel in the difference and the Bellman error is small regardless of how large the individual errors are.111111111In the extreme case, shifting all Q-values by the constant $c/(1-\gamma)$ leaves the Bellman error unchanged at zero while increasing value error by $c/(1-\gamma)$.
The second failure mode is specific to finite datasets: Fujimoto et al. (2022, Corollary 1) show that over an incomplete dataset, zero Bellman error is consistent with arbitrarily large value error, because the network can fit unobserved successor pairs to whatever values make the observed residuals vanish.112112112The Bellman equation uniquely identifies $Q^{\pi}$ when enforced over the entire MDP, but over an incomplete dataset it admits infinitely many solutions. Whenever a transition $(s^{\prime},a^{\prime})$ that is reachable from the dataset is not itself in the dataset, the network is free to set $Q(s^{\prime},a^{\prime})$ at unobserved pairs to whatever value makes the residual vanish on the observed ones, unconstrained by any loss. A network can thus reach near-zero training loss while the value function remains arbitrarily inaccurate over the full state space.
The dual failure mode appears in policy gradient methods: Ilyas et al. (2020) find that even when PPO’s surrogate objective improves monotonically, episode return can plateau or decline, because the surrogate gradient is poorly aligned with the gradient of the true return.113113113Ilyas et al. (2020) examine the surrogate objective used in Proximal Policy Optimization (Schulman et al., 2017). The PPO clipping mechanism is designed to keep policy updates within a trust region by bounding the probability ratio $\pi_{\theta}(a|s)/\pi_{\theta_{\text{old}}}(a|s)$. The gradient of the surrogate is poorly aligned with the gradient of the true return, particularly in later training phases where the policy has diverged from the behavior policy used to collect the replay data. The loss metric that practitioners monitor throughout training is measuring something other than what they care about.
@@ 6.2 The Reproducibility Crisis and Sensitivity to Random Seeds
Henderson et al. (2018) trained five leading policy gradient algorithms (PPO, TRPO, DDPG, TD3, SAC) on six MuJoCo benchmark environments, holding all hyperparameters fixed and varying only the random seed. The resulting learning curves from different seeds were non-overlapping: a seed that performed well under one algorithm performed comparably to a different algorithm’s best seeds, making cross-algorithm comparison unreliable.114114114Differences as large as 2,000 points in final episode return arose from seed variation alone. Agarwal et al. (2021b) quantify the damage: comparing point estimates from 5 runs per task on Atari 100k yields Type I error exceeding 50%, meaning a random noise injection appears beneficial in half of all comparisons.
Agarwal et al. (2021b) propose the interquartile mean (IQM) as a replacement for mean and median when comparing algorithms. The IQM discards the top and bottom 25% of runs before averaging, reducing sensitivity to outlier seeds. Using these tools and stratified bootstrap confidence intervals, Agarwal et al. (2021b) find that several widely-cited algorithmic improvements on Atari 100k vanish or reverse when statistical uncertainty is accounted for.115115115They also introduce performance profiles, which plot the fraction of tasks and seeds where an algorithm achieves performance above a threshold $\tau$, as $\tau$ varies from 0 to the maximum. Performance profiles reveal the full shape of the score distribution rather than collapsing it to a single statistic.116116116The fragility extends to hyperparameters. Eimer et al. (2023) conduct a systematic study of hyperparameter sensitivity across 6 algorithms and 17 environments, finding that default hyperparameters from published papers perform competitively in the specific environments used in those papers but generalize poorly across environments. Patterson et al. (2024) synthesize these findings into an empirical design handbook, recommending at minimum 10 seeds per configuration, IQM-based comparisons, and preregistration of hyperparameter search protocols.
@@ 6.3 Value Overestimation and Spikes
Fujimoto et al. (2018) observe that Double DQN’s correction is incomplete in continuous-action settings, where the target network and online network remain correlated through shared updates. They propose Clipped Double Q-learning: compute two Q-value estimates $Q_{1},Q_{2}$ with separate networks trained on the same data, and use $y=r+\gamma\min(Q_{1}(s^{\prime},a^{\prime}),Q_{2}(s^{\prime},a^{\prime}))$ as the Bellman target. The minimum operator introduces pessimistic underestimation, which is conservative but avoids the explosive positive bias.117117117Ciosek et al. (2019) note that clipped double Q can cause excessive pessimism under high uncertainty, proposing optimistic actor-critic as a counterweight.
@@
DQN prevents outright divergence, but van Hasselt et al. (2018) find that soft divergence, defined as a temporary spike in value estimates by more than 10% followed by recovery, occurs in the majority of DQN training runs across 57 Atari games.118118118Larger networks diverge more frequently than smaller ones, counter to the usual intuition that more expressive models should generalize better. The deadly triad does not announce itself as a training failure: the value estimates may spike and recover, leaving no visible trace in the loss curve while corrupting the policy.
Kumar et al. (2021) describe a continuous manifestation of the deadly triad: implicit under-parameterization.119119119Neural networks trained with bootstrapped TD objectives progressively lose their effective rank, with fewer and fewer neurons contributing distinct directions in the representation. This rank collapse is silent (training loss continues to decrease) but the network’s ability to represent new information degrades over time, a form of capacity loss distinct from but related to plasticity loss.
@@ 6.4 Plasticity Loss and Primacy Bias
Nikishin et al. (2022) identify primacy bias as a specific cause. Because the replay buffer is filled incrementally, early transitions are oversampled relative to later ones, and the network over-fits to early environment transitions, corrupting representations throughout the remainder of training.120120120Nikishin et al. (2022) show that 100 initial “priming” steps of excessive gradient updates degrade a SAC agent’s performance for hundreds of thousands of subsequent steps. Sokar et al. (2023) measure the fraction of dormant neurons (units with near-zero activation across the replay buffer) accumulating monotonically during training. Dohare et al. (2024) find that standard deep networks lose all plasticity within a few million gradient steps in continual learning tasks.
The proposed remedies fall into three categories: periodic resets, continual backpropagation, and architectural interventions.121121121Periodic resets reinitialize the last layers of the network while retaining the replay buffer, allowing the agent to forget overfit representations without discarding experience (Nikishin et al., 2022; D’Oro et al., 2023). Continual backpropagation replaces neurons with near-zero utility at each gradient step rather than waiting for a global reset (Dohare et al., 2024). Architectural interventions—layer normalization (Lyle et al., 2025), orthogonal initialization, spectral normalization—reduce the rate at which plasticity is lost by stabilizing gradient magnitudes and preventing weight norm growth.